Find F when c =37
c=5/9(f-32)

Answer by checkley71(8403)

37=5/9(F-32) OR 37(9)=5F-160 OR 333=5F-160 OR 5F=493 OR F=493/5 OR F=98.6
PROOF 37=5/9(98.6-32) OR 333=5(66.6) OR 333=333